OpenStudy (anonymous):

For which value is the rational expression x-6/5x-4 undefined?

OpenStudy (australopithecus):

well Remember you cannot divide by zero

OpenStudy (australopithecus):

since this is a fraction what values of x make the denominator zero

OpenStudy (australopithecus):

If you know that you know where the function is undefined

OpenStudy (australopithecus):

so set 5x-4 = 0 and solve for x

OpenStudy (australopithecus):

5x = 4 x = 4/5

OpenStudy (australopithecus):

thus x=/= 4/5

OpenStudy (australopithecus):

In set notation it would be (-infinity, 4/5) U (4/5, +infinity)

OpenStudy (anonymous):

I see what you did there, but this is what I got,

OpenStudy (anonymous):

x/5x-4 - 6/5x-4

OpenStudy (anonymous):

\[5x+\sqrt{4}\]

OpenStudy (anonymous):

\[5x-\sqrt{4}\]

OpenStudy (australopithecus):

I have no idea what you are doing link23

OpenStudy (anonymous):

x= -3 and x= 7

OpenStudy (anonymous):

then I got x-6/(x-3)(x+7) = x=5/4

OpenStudy (australopithecus):

well lets put x=-3, and x=7 into the equation x-6/5x-4 ((-3) - 6)/(5(-3) - 4) = -9/-19 x = -3 is clearly in the function

OpenStudy (australopithecus):

link23 you are over complicating it.

OpenStudy (australopithecus):

You are just checking what x values cannot be subbed into the equation. Since you cannot divide by zero you simply have to see what values of x make the denominator zero. No manipulation of simplification of the fraction as a whole is needed in this case

OpenStudy (australopithecus):

*No manipulation or simplification of the fraction as a whole is needed in this case

OpenStudy (australopithecus):

furthermore x-6/(x-3)(x+7) =/= x-6/5x-4 as x-6/(x-3)(x+7) = x-6/(x^(2) + 4x - 21)

OpenStudy (australopithecus):

I dont know if you are trolling me or not, the question has been answered if you are still having problems with this consult your teacher/prof.